OpenStudy (anonymous):

From 52 cards, 2 cards are drawn. What is the probability that both are of the same suit?

OpenStudy (anonymous):

\[\frac{4\binom{13}{2}}{\binom{52}{2}}\]

OpenStudy (anonymous):

what does the numbers mean on top of each other? my teacher uses P for permutation and C for combination

OpenStudy (anonymous):

OK ! \[\frac{4C_{13}^2}{C_{52}^2}\]

OpenStudy (anonymous):

i'm sorry i don't understand the notation. i write it like, 52C2 or something like that... i dont get what the numbers on top of each other mean

OpenStudy (anonymous):

OK ! \[\frac{4\times(13C2)}{52C2}\]

OpenStudy (anonymous):

thank you!